OpenStudy (anonymous):

derivative of 3e^3 and e^x/3 is what?

OpenStudy (anonymous):

first one is a number. second one is a function of x. what do you mean by "derivation"?

OpenStudy (anonymous):

f(x)=(e^x/3)-(3e^3)

OpenStudy (anonymous):

question is...

OpenStudy (anonymous):

their respective derivates will be wat?

OpenStudy (anonymous):

Thank you to clarify it, Satellite73 :)

OpenStudy (anonymous):

if it is "find the derivative" the answer is \[f'(x)=\frac{1}{3}e^{\frac{x}{3}}\] by the chain rule

OpenStudy (anonymous):

cos i know the e is a constant,and when taking derivates u ignore the constant,but the example we did in class is different form this

OpenStudy (anonymous):

not with chain rule

OpenStudy (anonymous):

(1/3) e^(x/3)

OpenStudy (anonymous):

i know that the derivative of 3x^3 is 9x^2

OpenStudy (anonymous):

The 2nd one is constant, so derivative = 0

OpenStudy (anonymous):

thanks alot,u got my question

OpenStudy (anonymous):

straight forward formula: (e^u) ' = u' e^u

OpenStudy (anonymous):

thank you green one...those e^u problem continue to plague me